OpenStudy (anonymous):

Coffee-cup calorimeter When a 100.0 mL of 0.200 M NaOH solution is mixed with a 100.0 mL of 0.200 M HCl solution in a coffee-cup calorimeter, the temperature of the solution increases from 25.00 oC to 26.35 oC. Assuming that the specific heat of the solution is 4.184 J/(g oC) and that the calorimeter itself absorbs a negligible amount of heat, calculate the heat of the reaction (in kJ/mol) for the neutralization reaction.

OpenStudy (aaronq):

q=mCdT solve for q did they give you densities of the solutions?

OpenStudy (anonymous):

There is not density of the solutions. I learned to do q=mct with no d. There is 1g/ml density of the water. Which i think is absorbing the heat. Perphaps the heat gained by the water= heat lost by solutions. I also used the forumla H=U+NRT. The anwser is supposed to be -56 kj/mole but how do i get this?

OpenStudy (aaronq):

dT = delta T = (T2-T1)= change in temperature

OpenStudy (aaronq):

try using the density of water for the solution, 200mL